If 1 is added to five times a number, the result is equal to 6 more then four times the number. Find the number
Answer by 4419875(21)

Assign:
x - the number
So by translating we say that five times(5*x) a number:(5x+1) is equal(=) to six more then(+6) four times(4*x) the number:(4x+6).
Then, we organize the equation:
5x+1 = 4x+6
Isolate x by applying subtraction property:
5x+1-4x=4x+6-4x
x+1=6
Apply again the property
x+1-1=6-1
x=5
Let's check.
5(5)+1 = 4(5)+6
25+1 = 20+6
26 = 26
Therefore our answer is correct that our number(x) is 5